### v3

# Noise Sensitivity and Learning Lower Bounds for Hierarchical Functions

Abstract.
Recent works explore deep learning’s success by examining functions or data with hierarchical structure. To study the learning complexity of functions with hierarchical structure, we study the noise stability of functions with tree hierarchical structure on independent inputs. We show that if each function in the hierarchy is $\varepsilon$-far from linear, the noise stability is exponentially small in the depth of the hierarchy.
Our results have immediate applications for agnostic learning. In the Boolean setting using the results of Dachman-Soled, Feldman, Tan, Wan and Wimmer (2014), our results provide Statistical Query super-polynomial lower bounds for agnostically learning classes that are based on hierarchical functions.
We also derive similar SQ lower bounds based on the indicators of crossing events in critical site percolation. These crossing events are not formally hierarchical as we define but still have some hierarchical features as studied in mathematical physics.
Using the results of Abbe, Bengio, Cornacchiam, Kleinberg, Lotfi, Raghu and Zhang (2022), our results imply sample complexity lower bounds for learning hierarchical functions with gradient descent on fully connected neural networks.
Finally in the Gaussian setting, using the results of Diakonikolas, Kane, Pittas and Zarifis (2021), our results provide super-polynomial lower bounds for agnostic SQ learning.
1. Introduction
Our goal is to bridge two lines of research in learning theory. The first line proposed that the success of neural networks can be explained in terms of hierarchical structures in the data. The second line of work provides learning lower bounds for noise sensitive functions, or classes of such functions.
Multiple works have argued that the success of deep learning is intimately related to hierarchical properties in data generative processes. This hypothesis has been independently proposed by multiple researchers, including:
  • Bruna and Mallat’s wavelet scattering networks [6],
  • Mhaskar, Liao, and Poggio’s compositional function models [19],
  • Patel, Nguyen, and Baraniuk’s hierarchical rendering models [24].
Mossel [22] provided a rigorous attempt to establish computational hardness, though subsequent work [18, 16] only demonstrated low-degree hardness. In terms of compositions of functions, Telgarsky [30] presents a notable result demonstrating an iteratively composed simple function that can be precisely expressed as a deep narrow network, yet requires super-polynomial computational complexity for small network depth.
The second research line establishes learning lower bounds in terms of the Fourier expansion of functions or classes of functions. This includes in particular SQ lower bounds in terms of the noise-sensitivity in both the Boolean case [8] and the Gaussian case [9] as well as sample complexity lower bounds for agnostically learning such functions using fully connected neural networks and stochastic gradient descent [1].
In our main result we show that hierarchical functions, i.e., functions obtained by decompositions of non-linear functions on disjoint independent inputs, are noise sensitive. This implies SQ lower bounds for agnostically learning such functions in both the Boolean and the Gaussian case as well as sample lower bounds for learning such functions using neural networks.
Our work builds upon a long line of work in discrete Fourier analysis. The fact that recursively composing the majority function on $3$ bits leads to a noise sensitive function is folklore in the area, see e.g., [4, 5]. Mossel and O’Donnell [20] further study decompositions of monotone functions and construct such decompositions that are essentially “most” noise sensitive among monotone functions. We also note that recursive constructions, i.e., Sipser functions, provide the best results in circuit lower bounds for circuits with AND, OR and NOT gates [14].
The main contribution of the current paper is by showing that noise sensitivity of composition of non-linear functions on non-overlapping inputs is generic—the functions to be composed do not have to be identical, to be chosen carefully, or to be balanced or have any other property other than being somewhat uncorrelated with linear functions of their inputs. Our results hold for all product measures, both discrete and continuous.
1.1. Our Results
1.1.1. Multilinear functions
To begin with we consider the case where all functions involved are multilinear. This includes in particular the case where all functions have binary inputs and outputs which is the classical case where analysis of Boolean functions and noise sensitivity are discussed. In this case we obtain the following result concerning a single multilinear function. Precise definitions are provided in Section 2, though we comment that $d(f,\operatorname{Lin})$ quantifies the non-linearity of $f$.
Lemma 1.1.
Suppose $\left\{(X_{i},Y_{i}):i\in[n]\right\}$ are mutually independent, where for some $0\leq\rho\leq 1$, we have $\operatorname{Corr}(X_{i},Y_{i})\leq\rho$ for all $i$. Suppose $X_{i}$ and $Y_{i}$ have the same first and second moments, i.e., $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]$ and $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]<\infty$, for all $i\in[n]$. If $f:\mathbb{R}^{n}\to\mathbb{R}$ is a multilinear function satisfying $d(f,\operatorname{Lin})\geq\varepsilon>0$, then
(1.1) $\operatorname{Corr}(f(X),f(Y))\leq(1-\varepsilon)\rho+\varepsilon\rho^{2}.$
Remark 1.2.
Such a $Y$ can always be generated by setting $Y_{i}=X_{i}$ with probability $\rho$ and otherwise letting $Y_{i}$ be independent and identically distributed as $X_{i}$, independently for all $i$; note that this is not in general the unique construction. Moreover, this construction yields $X$ and $Y$ that have the same marginal distribution, which does not need to be true in general for Lemma 1.1 to hold.
Applying Lemma 1.1 inductively yields the following result demonstrating exponential decay in noise stability for a hierarchical multilinear function, precisely defined in Section 2.3.
Theorem 1.3.
Suppose $\left\{(X_{i},Y_{i}):i\in[n]\right\}$ are mutually independent, where for some $0\leq\rho<1$, we have $\operatorname{Corr}(X_{i},Y_{i})\leq\rho$ for all $i$. Suppose $X_{i}$ and $Y_{i}$ have the same first and second moments, i.e., $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]$ and $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]<\infty$, for all $i\in[n]$. If $f:\mathbb{R}^{n}\to\mathbb{R}$ is a hierarchical multilinear function of non-linearity $\varepsilon>0$ and depth $d$, then for any $\delta<\varepsilon$,
$\operatorname{Corr}(f(X),f(Y))\leq\left(1-\delta\right)^{d-C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)}$
for some constant $C_{\varepsilon,\delta}>0$ depending only on $\varepsilon$ and $\delta$.
As discussed in Remark 2.2, this bound is essentially tight, as it is impossible to have exponential rate smaller than $1-\varepsilon$. Remark 2.3 discusses how this can be improved to doubly exponential decay in depth in the case where each component function has 0 correlation with linear functions, such as $t$-resilient functions from cryptography, which are further discussed in Remark 2.4, and the parity bit function as further discussed in Remark 2.5.
1.1.2. General functions
The results above leave something to be desired, as neural network activation functions, notably ReLU, are typically not multilinear in their inputs. However, some care is needed in this general case as the following examples show.
Example 1.4.
Consider a hierarchical function with a binary tree structure, where at even levels we have component functions $f(x_{1},x_{2})=\cos(\pi x_{1})$ and at odd levels $g(x_{1},x_{2})=\frac{\arccos(x_{1})}{\pi}$, where each of the component variables takes values in $[0,1]$. For example, such a function of depth 2 would be given by $g(f(x_{1},x_{2}),f(x_{3},x_{4}))$. We note that $f$ and $g$ are clearly non-linear, but any such function of even depth will output $x_{1}$, which is as noise stable as possible, and there is no decay in correlation as depth increases.
Example 1.5.
The phenomenon above does not require the functions to have continuous inputs. In fact, it suffices for the inputs and outputs of the functions to take on a small number of values, e.g., four. For example, we can consider the input distributions to be uniform over $\left\{-1,1\right\}^{3^{d}}$, i.e., $3^{d}$ i.i.d. variables uniformly distributed on $\{-1,1\}$. In our correlated setup for noise sensitivity, we consider two such random vectors $X=(X_{1},\dots,X_{n})$ and $Y=(Y_{1},\dots,Y_{n})$ coupled such that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are mutually independent, and $\operatorname{Corr}(X_{i},Y_{i})=\rho$ for all $i$.
In our construction, our functions can output at most four values, $\pm 10\pm 1$, and our overall hierarchical function has a ternary tree structure. For a number $x=10b_{1}+b_{2}$ for $b_{1},b_{2}\in\{-1,1\}$, let us write $B_{1}(x)=b_{1}$ and $B_{2}(x)=b_{2}$. We can then take the functions at the lowest layer (depth $d$) to be
$f(x_{1},x_{2},x_{3})=x_{1}+10\operatorname{maj}(x_{1},x_{2},x_{3}),$
where $\operatorname{maj}(x_{1},x_{2},x_{3})=2\cdot 1_{\left\{x_{1}+x_{2}+x_{3}>0\right\}}-1$ is the majority function, i.e., takes value 1 if a majority of inputs take value 1, and takes value 1-1 if a majority of inputs takes value 1-1. We let the remaining functions be
$g(x_{1},x_{2},x_{3})=B_{2}(x_{1})+10\operatorname{maj}(B_{1}(x_{1}),B_{1}(x_{2}),B_{1}(x_{3})),$
It is easy to see that all functions involved are far from linear, but the hierarchical function on $3^{d}$ inputs computes the first input plus $10$ times the recursive majority of all $3^{d}$ inputs, and is therefore very noise stable as it has constant correlation with $x_{1}$.
To account for potential non-linear relationships, we strengthen our notions of correlation in both our assumptions and our conclusions. The precise definitions are provided in Section 3, but loosely speaking, $\operatorname{Corr}(L^{2}(X),L^{2}(Y))$ denotes the maximum correlation between any two functions where one is a function of $X$ and the other is a function of $Y$. The analog of Lemma 1.1 in this more general setting is the following lemma.
Lemma 1.6.
Suppose $\left\{(X_{i},Y_{i}):i\in[n]\right\}$ are mutually independent, where for some $0\leq\rho\leq 1$, we have $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))\leq\rho$ for all $i$. If $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfies $\mathbb{E}[f(X)^{2}],\mathbb{E}[f(Y)^{2}]<\infty$ and
(1.2) $\operatorname{Corr}(L^{2}(f(X)),L^{2}(X_{1})+\cdots+L^{2}(X_{n})),\operatorname{Corr}(L^{2}(f(Y)),L^{2}(Y_{1})+\cdots+L^{2}(Y_{n}))\leq\sqrt{1-\varepsilon}$
for some $\varepsilon>0$, then
$\operatorname{Corr}(L^{2}(f(X)),L^{2}(f(Y)))\leq(1-\varepsilon)\rho+\varepsilon\rho^{2}.$
In Section 3.3, we justify why (1.2) is the natural analog of the condition $d(f,\operatorname{Lin})\geq\varepsilon$ in Lemma 1.1.
Remark 1.7.
It is straightforward to check that the construction in Remark 1.2 still satisfies the stronger condition $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))=\rho$: letting $Z_{i}$ be an independent copy of $X_{i}$ such that $Y_{i}=X_{i}$ with probability $\rho$ and $Y_{i}=Z_{i}$ otherwise, we have, for $f\in L^{2}(X_{i})$ and $g\in L^{2}(Y_{i})$ normalized so that $\mathbb{E}[f]=\mathbb{E}[g]=0$ and $\mathbb{E}[f^{2}]=\mathbb{E}[g^{2}]=1$, by the Cauchy–Schwarz inequality
$\operatorname{Corr}(f(X_{i}),g(Y_{i}))=\rho\mathbb{E}[f(X_{i})g(X_{i})]+(1-\rho)\mathbb{E}[f(X_{i})g(Z_{i})]\leq\rho\cdot 1+(1-\rho)\cdot 0=\rho,$
where equality is achieved when $f=g$ as functions $\mathbb{R}\to\mathbb{R}$.
Using Lemma 1.6, we obtain the following general version of Theorem 1.3.
Theorem 1.8.
Suppose $\left\{(X_{i},Y_{i}):i\in[n]\right\}$ are mutually independent, where for some $0\leq\rho<1$, we have $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))\leq\rho$ for all $i$. If $f:\mathbb{R}^{n}\to\mathbb{R}$ is a hierarchical function of depth $d$ and non-separability $\varepsilon>0$ in both $L^{2}(X)$ and $L^{2}(Y)$, then for any $\delta<\varepsilon$,
$\operatorname{Corr}(L^{2}(f(X)),L^{2}(f(Y)))\leq(1-\delta)^{d-C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)}$
for some constant $C_{\varepsilon,\delta}>0$ depending only on $\varepsilon$ and $\delta$.
1.2. Applications
Theorem 1.3 immediately implies a low-degree bound on a hierarchical multilinear function—in particular we get that the correlation $M$ of such a function $f$ with a degree-$D$ polynomial is bounded by
(1.3) $M\leq\rho^{-D/2}(1-\delta)^{d/2-C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)},$
for any $\rho\in(0,1)$; see Lemma 2.6. Taking $\rho=1-e^{-cd}$ for some small $c\in\left(0,\frac{1}{2C_{\varepsilon,\delta}}\right)$, we get a degree bound that is exponential in the circuit depth, i.e., in order to get $M=\Omega(1)$ correlation, one requires $D=e^{\Omega(d)}$.
In the Boolean setting, which is contained in our framework of multilinear functions, using the results of Dachman-Soled, Feldman, Tan, Wan and Wimmer [8], the bound (1.3) implies super-polynomial lower bounds for Statistical Query (SQ) agnostic learning of concept classes based on hierarchical multilinear functions. We refer readers to [8] for background on agnostic learning in the SQ model.
Theorem 1.9.
For any constant $\varepsilon>0$, for any hierarchical multilinear function $f:\left\{-1,1\right\}^{n^{1/3}}\to[-1,1]$ of depth $d=\Theta(\log n)$ and non-linearity $\varepsilon$, let $\mathcal{C}$ be the concept class of functions $g:\left\{-1,1\right\}^{n}\to[-1,1]$ for which there exist distinct $i_{1},\dots,i_{n^{1/3}}\in[n]$ such that $g(x)=f(x_{i_{1}},\dots,x_{i_{n^{1/3}}})$ for all $x$. Then any SQ algorithm for agnostically learning $\mathcal{C}$ with excess error of at most $\frac{1}{2}-o(1)$ has complexity at least $n^{\Omega\left(\frac{\log n}{\log\log(n)}\right)}$.
It is natural to expect that similar results hold for other types of hierarchical functions, not necessarily as defined thus far. A classical example of such functions is crossing events in critical percolation; we refer readers to [13] for background on percolation. Renormalization group arguments in statistical physics (see [3] for an overview of renormalization groups and [29] for a more focused treatment on percolation) have argued that such events are self-similar in the sense that the event views at different scales all look the same. Thus, like the example of recursive majority of $3$, crossing events can be summarized (in some sense) over squares to obtain again a picture of critical percolation. While this picture has not been proven in a rigorous way, the breakthrough results in statistical physics by Schramm, Werner, Smirnov and others (see, e.g., [26, 27, 28]) allowed mathematicians to prove related detailed estimates about these functions and their Fourier spectra, which suffices for our noise sensitivity estimates. In particular, applying a similar proof strategy as that of Theorem 1.9 to previously known results by Garban, Pete, and Schramm [11] on the noise sensitivity of critical site percolation on the triangular grid, we show percolation crossing events also create a concept class of functions that imply similar super-polynomial SQ lower bounds for agnostic learning.
Theorem 1.10.
Let $f$ be the $\pm 1$ indicator function of the left-to-right crossing event of the square $0,n]^{2}$ in critical site percolation on the triangular grid. Let $N=\Theta(n^{2})$ be the number of sites in this square, so that $f$ is a function $\left\{-1,1\right\}^{N}\to\left\{-1,1\right\}$. Let $\mathcal{C}$ be the concept class of functions $g:\left\{-1,1\right\}^{N^{3}}\to\left\{-1,1\right\}$ for which there exist distinct $i_{1},\dots,i_{N}\in\left[N^{3}\right]$ such that $g(x)=f(x_{i_{1}},\dots,x_{i_{N}})$ for all $x$. Then any SQ algorithm for agnostically learning $\mathcal{C}$ with excess error of at most $\frac{1}{2}-o(1)$ has complexity at least $n^{\Omega\left(\frac{\log n}{\log\log(n)}\right)}$.
Analogously, Theorem 1.8 immediately implies a low-degree bound on hierarchical functions, similar to (1.3). The correlation $M$ of such a function $f$ with a function of Efron–Stein degree at most $D$ (see Section 3.2 for precise definitions), i.e., a sum of functions that each depend on at most $D$ of the $n$ variables, is bounded by
(1.4) $M\leq\rho^{-D/2}(1-\delta)^{d/2-C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)};$
see Lemma 3.4. As before, this implies a degree bound exponential in depth.
When our inputs are Gaussian, using the results of Diakonikolas, Kane, Pittas and Zarifis [9], Theorem 1.8 implies super-exponential lower bounds for agnostic learning in the SQ model for concept classes based on hierarchical functions.
Theorem 1.11.
Fix an arbitrary constant $\varepsilon>0$. Let positive integers $n$ and $m$ satisfy $m\leq n^{\alpha}$ for any constant $\alpha<\frac{1}{2}$. For any hierarchical multilinear function $f:\mathbb{R}^{m}\to\left\{\pm 1\right\}$ of depth $d=\Theta(\log n)$ and non-separability $\varepsilon$ in $L^{2}(X)$, where $X\sim\mathcal{N}(0,I_{m})$ is a $m$-dimensional spherical Gaussian of unit variance, let $\mathcal{C}$ be the concept class consisting of the functions of the form $F(x)=f(Px)$, where $P\in\mathbb{R}^{m\times n}$ is any matrix satisfying $PP^{T}=I_{m}$. Then there exists constant $c=c_{\varepsilon}>0$ such that any SQ learning algorithm that agnostically learns $\mathcal{C}$ over $\mathcal{N}(0,I_{n})$ to $L^{2}$-error $\operatorname{OPT}+\frac{1}{8}-o(1)$ either requires queries with tolerance at most $n^{-\Omega(n^{c})}$ or makes at least $2^{n^{\Omega(1)}}$ queries.
Our final application is for learning such functions using stochastic gradient descent on fully connected deep neural networks. The results of [1] which build on [2] imply that for functions as above, in their setting of fully connected neural networks and their initialization, the complexity of learning the function using their version of gradient descent is proportional exponential in the depth of the circuit. In the special case where all functions are resilient, i.e., have 0 correlation with linear function, the lower bound becomes doubly exponential in the depth, i.e., exponential in the input size. More formally, [1, Theorem 1] immediately implies the following.
Corollary 1.12.
Let $f:\{-1,1\}^{n}\to\{-1,1\}$ be a hierarchical multilinear function of non-linearity $\varepsilon>0$ and depth $d=\Omega(\log n)$ that is balanced, i.e., $\mathbb{E}[f(X)]=0$ where $X\sim\operatorname{Unif}(\left\{-1,1\right\}^{n})$. Let $\bar{f}$ be the $2n$-extension of $f$, given by $\bar{f}(x_{1},\ldots,x_{2n})=f(x_{1},\ldots,x_{n})$. Consider a fully connected neural network of size $E$ with initialization that is target agnostic in the sense that it is invariant under permutations of the inputs. Let $f^{(t)}_{\mathrm{NN}}$ denote the output of noisy-GD with gradient range $A$, batch-size $b$, learning rate $\gamma$, and noise scale $\sigma$ (see [1, Definition 8]) after $t$ time steps. Then, if $1-\rho\leq 0.25$ and $b$ is sufficiently large, then for any $\delta<\varepsilon$ the generalization error is at least 0.5o(1)0.5-o(1) as long as
$\gamma t\sqrt{E}A(1-\delta)^{d-C_{\varepsilon,\delta}\log(\frac{1}{1-\rho})}\ll\sigma.$
In the case where $\varepsilon=1$, we get generalization error at least 0.5o(1)0.5-o(1) as long as $\gamma t\sqrt{E}A\rho^{2^{d-2}}\ll\sigma$.
Note that the bound in the case $\varepsilon<1$ is polynomial in $n$, while we get super-polynomial lower bound in the case where $\varepsilon=1$ (see Remark 2.3 below).
We conjecture that the lower bounds of Abbe, Bengio, Cornacchiam, Kleinberg, Lotfi, Raghu and Zhang [1] should extend in a straightforward way to other product spaces.
We finally note briefly a different set of results involving noise sensitivity and deep neural networks, i.e., the results of [17], which studied the noise sensitivity of ReLU networks with i.i.d. Gaussian weights.
1.3. Section Overview
In Section 2, we introduce the necessary definitions to prove the results concerning multilinear functions. Then in Section 3, we first introduce the necessary formalization of correlated probability spaces in Section 3.1, then introduce the Efron–Stein decomposition in Section 3.2 and its fundamental properties, especially in relation to Markov operators; finally, properly equipped, we can prove the results concerning general functions.
2. Noise stability of compositions of multilinear functions
2.1. Preliminary setup
For any probability space $(\Omega,\mathbb{P})$, function $f\in L^{2}(\Omega,\mathbb{P})$, and subset $V\subseteq L^{2}(\Omega,\mathbb{P})$, define the (squared) distance between $f$ and $V$ by
$d(f,V)=\inf_{g\in V}\frac{\mathbb{E}[(f-g)^{2}]}{\operatorname{Var}(f)}.$
If $f$ is almost surely constant, i.e., $\operatorname{Var}(f)=0$, we define $d(f,V)=0$. Similarly, throughout this paper we use the convention $\operatorname{Corr}(X,Y)=0$ if either $X$ or $Y$ are almost surely constant. We now specialize to the case where $\mathbb{P}$ is a product measure on product space $\Omega$ over laws of real-valued random variables $X_{1},X_{2},\dots,X_{n}$. In the product measure, these random variables are mutually independent, and we additionally assume the random variables have finite second moment, i.e., $\mathbb{E}[X_{i}^{2}]<\infty$ for all $i\in[n]$. Beyond this, we make no further assumptions on the distributions of the random variables, e.g., some may be discrete, others continuous, or neither. We will consider the setting where $f\in L^{2}(\Omega,\mathbb{P})$ is multilinear, i.e., $f$ is of the form
$f(x_{1},\dots,x_{n})=\sum_{S\subseteq[n]}a_{S}\prod_{i\in S}x_{i}$
for some coefficients $a_{S}\in\mathbb{R}$. Note that any function of the above form is in $L^{2}(\Omega,\mathbb{P})$. In the Boolean case, i.e., when $X_{1},\dots,X_{n}$ are each random variables supported on at most two points, note that all $f\in L^{2}(\Omega,\mathbb{P})$ are multilinear in $X_{1},\dots,X_{n}$.
Let $\operatorname{Lin}$ denote the subspace of $L^{2}(\Omega,\mathbb{P})$ spanned by linear functions, i.e.,
$\operatorname{Lin}=\operatorname{span}(\{1,x_{1},x_{2},\dots,x_{n}\}).$
For each $i\in[n]$, define the normalized random variable $\widetilde{X}_{i}=\frac{X_{i}-\mathbb{E}[X_{i}]}{\sqrt{\operatorname{Var}(X_{i})}}$, which has mean 0 and variance 1, and similarly denote $\widetilde{x}_{i}=\frac{x_{i}-\mathbb{E}[X_{i}]}{\sqrt{\operatorname{Var}(X_{i})}}$ for all $x_{i}\in\mathbb{R}$. Then for any multilinear $f$, we let $\widehat{f}(S)\in\mathbb{R}$ for $S\subseteq[n]$ denote the unique coefficients satisfying
$f(x_{1},\dots,x_{n})=\sum_{S\subseteq[n]}\widehat{f}(S)\prod_{i\in S}\widetilde{x}_{i}.$
For each $S\subseteq[n]$, we let $\chi_{S}:\mathbb{R}^{n}\to\mathbb{R}$ denote the function given by
$\chi_{S}(x_{1},\dots,x_{n})=\prod_{i\in S}\widetilde{x}_{i}.$
By independence of $X_{1},\dots,X_{n}$, we have $\left\{\chi_{S}:S\subseteq[n]\right\}$ is a set of orthonormal multilinear polynomials: for distinct $S,S^{\prime}\subseteq[n]$, we have $\mathbb{E}[\chi_{S}\chi_{S^{\prime}}]=0$ and $\mathbb{E}[\chi_{S}^{2}]=1$. Also note that $\mathbb{E}[\chi_{S}]=1_{\left\{S=\emptyset\right\}}$. Hence, we have
$\mathbb{E}[f^{2}]=\sum_{S\subseteq[n]}\widehat{f}(S)^{2}\quad\text{and}\quad\operatorname{Var}(f)=\sum_{S\neq\emptyset}\widehat{f}(S)^{2}.$
From this, it is easy to see that
$d(f,\operatorname{Lin})=\frac{\displaystyle\sum_{\left\lvert S\right\rvert\geq 2}\widehat{f}(S)^{2}}{\displaystyle\sum_{S\neq\emptyset}\widehat{f}(S)^{2}}.$
For any $S\subseteq[n]$, we let $X_{S}$ denote the random vector consisting of the $X_{i}$ for $i\in S$, and similarly define $Y_{S}$, as well as $x_{S}$ and $y_{S}$ for $x,y\in\mathbb{R}^{n}$. For example, $\chi_{S}(x)$ depends only on $x_{S}$.
2.2. Main lemma
Lemma 1.1 shows that if we have some noisy observation $Y_{i}$ of each $X_{i}$, namely $\operatorname{Corr}(X_{i},Y_{i})\leq\rho$ for all $i\in[n]$ for some constant $\rho<1$, and $f$ is far from linear, then $\operatorname{Corr}(f(X),f(Y))<\rho$, with this correlation displaying exponential decay with respect to application of $f$. We now prove this lemma.
Proof of Lemma 1.1.
Note that because $X_{1},\dots,X_{n}$ are mutually independent, and $Y_{1},\dots,Y_{n}$ are also mutually independent, we have $\operatorname{Var}(f(X))=\operatorname{Var}(f(Y))$ as
$\mathbb{E}[\chi_{S}(X)\chi_{S^{\prime}}(X)]=\mathbb{E}[\chi_{S}(Y)\chi_{S^{\prime}}(Y)]=1_{\left\{S=S^{\prime}\right\}}.$
We similarly also have $\mathbb{E}[f(X)]=\mathbb{E}[f(Y)]=\widehat{f}(0)$ as $\mathbb{E}[\chi_{S}(X)]=\mathbb{E}[\chi_{S}(Y)]=1_{\left\{S=\emptyset\right\}}$. Thus, without loss of generality, we can rescale and shift $f$ so that $\operatorname{Var}(f)=1$ and $\mathbb{E}[f]=0$. Then we expand
$\operatorname{Corr}(f(X),f(Y))$ $=\mathbb{E}[f(X)f(Y)]=\sum_{S\neq\emptyset}\sum_{S^{\prime}\neq\emptyset}\widehat{f}(S)\widehat{f}(S^{\prime})\mathbb{E}[\chi_{S}(X)\chi_{S^{\prime}}(Y)].$
Note that if $S\neq S^{\prime}$, we have $\mathbb{E}[\chi_{S}(X)\chi_{S^{\prime}}(Y)]=0$. Then either $S\not\subseteq S^{\prime}$ or $S^{\prime}\not\subseteq S$. Say $S\not\subseteq S^{\prime}$. Then
$\mathbb{E}[\chi_{S}(X)\chi_{S^{\prime}}(Y)]$ $=\mathbb{E}[\mathbb{E}[\chi_{S}(X)\chi_{S^{\prime}}(Y)|(X_{S^{\prime}},Y_{S^{\prime}})]]$
$=\mathbb{E}[\mathbb{E}[\chi_{S}(X)|(X_{S^{\prime}},Y_{S^{\prime}})]\chi_{S^{\prime}}(Y)]$
$=\mathbb{E}[0\cdot\chi_{S^{\prime}}(Y)]$
$=0.$
The case $S^{\prime}\not\subseteq S$ follows symmetrically. If $S=S^{\prime}$, then by independence of the $(X_{i},Y_{i})$ we have
$\mathbb{E}[\chi_{S}(X)\chi_{S}(Y)]$ $=\prod_{i\in S}\mathbb{E}[\widetilde{X}_{i}\widetilde{Y}_{i}]=\prod_{i\in S}\operatorname{Corr}(X_{i},Y_{i})\leq\rho^{\left\lvert S\right\rvert}$.
Thus,
$\operatorname{Corr}(f(X),f(Y))\leq\sum_{S\neq\emptyset}\rho^{\left\lvert S\right\rvert}\widehat{f}(S)^{2}.$
Recalling that
$\operatorname{Var}(f)=\sum_{S\neq\emptyset}\widehat{f}(S)^{2}=1$
while
$d(f,\operatorname{Lin})=\sum_{\left\lvert S\right\rvert\geq 2}\widehat{f}(S)^{2}\geq\varepsilon,$
we obtain the bound
$\operatorname{Corr}(f(X),f(Y))\leq(1-\varepsilon)\rho+\varepsilon\rho^{2},$
as desired. ∎
Note that this bound is sharp when the conditions in the statement of the lemma are tight, i.e., $\operatorname{Corr}(X_{i},Y_{i})=\rho$ for all $i$, and $d(f,\operatorname{Lin})=\varepsilon$, and $f$ is of degree 2.
Remark 2.1.
In the Boolean setting, i.e., when the $X_{i}$ and $Y_{i}$ are all Boolean random variables, when $\operatorname{Corr}(X_{i},Y_{i})=\rho$ for all $i$, then the quantity $\operatorname{Corr}(f(X),f(Y))$ that we study in Lemma 1.1 coincides with the well-studied noise stability $\operatorname{Stab}_{\rho}(f)$. We refer readers to [23] for a comprehensive overview of the analysis of Boolean functions. As we wish to state results that apply for general settings beyond just the Boolean setting, we will not use the specific definition of noise stability $\operatorname{Stab}_{\rho}(f)$, though the spirit remains the same.
2.3. Hierarchical multilinear functions
We now consider functions consisting of a composition of multilinear functions in a hierarchical structure. Specifically, for any positive integers $n$ and $d$, we define a function $f:\mathbb{R}^{n}\to\mathbb{R}$ to be a hierarchical multilinear function of non-linearity $\varepsilon>0$ and depth $d$ recursively as follows. A hierarchical multilinear function of non-linearity $\varepsilon>0$ and depth 1 is simply a function satisfying the conditions of Lemma 1.1. Note that $d(f,\operatorname{Lin})$ does not depend on the underlying random variables $X_{1},\dots,X_{n}$. A hierarchical multilinear function of non-linearity $\varepsilon>0$ and depth $d\geq 2$ is a function $f:\mathbb{R}^{n}\to\mathbb{R}$ of the form
$f(x)=g(h_{1}(x_{\mathcal{P}_{1}}),\dots,h_{m}(x_{\mathcal{P}_{m}}))$
for some partition $\mathcal{P}$ of $[n]$ with $m\geq 2$ parts and hierarchical multilinear functions $g,h_{1},\dots,h_{m}$ of non-linearity $\varepsilon$, where $g$ has depth 1 and each $h_{i}$ has depth at least d1d-1. While this definition allows for the $h_{i}$ to have depth greater than or equal to $d$, note that any depth-$d$ hierarchical multilinear function of non-linearity $\varepsilon>0$ must be a function of at least $2^{d}$ variables, so our recursive definition does terminate and is well-defined.
As previously mentioned, all functions on Boolean random variables are multilinear, so any function on Boolean random variables of this hierarchical form whose component functions all have Boolean outputs, i.e., take on two values, except for potentially the top component function, are hierarchical multilinear functions. Thus, our definition of a hierarchical multilinear function includes all hierarchical Boolean functions.
We can now prove Theorem 1.3, which shows that $\operatorname{Corr}(f(X),f(Y))$ decays exponentially in depth $d$, for hierarchical multilinear functions of positive non-linearity.
Proof of Theorem 1.3.
Let $g(x)=(1-\varepsilon)x+\varepsilon x^{2}$. For positive integer $m\geq 1$, let $g^{m}$ denote the composition of $g$ a total of $m$ times. Because $X_{i}$ and $Y_{i}$ have the same first and second moments, and because the $(X_{i},Y_{i})$ are mutually independent across $i$, any multilinear function $h$ will have $\mathbb{E}[h(X)]=\mathbb{E}[h(Y)]$ and $\mathbb{E}[h(X)^{2}]=\mathbb{E}[h(Y)^{2}]$, and thus we can repeatedly apply Lemma 1.1 for all component functions of our hierarchical multilinear function $f$, as the first and second moment conditions of Lemma 1.1 inductively hold. By repeatedly applying Lemma 1.1 throughout the hierarchical structure of $f$, and noting that $g$ is an increasing function on $[0,1]$ with $0\leq g(x)<x$ for all $0\leq x<1$, we have $\operatorname{Corr}(f(X),f(Y))\leq g^{d}(\rho)$, so it suffices to show $g^{d}(\rho)\leq\left(1-\delta\right)^{d-C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)}$ for all $d\geq 1$ and $0\leq\rho<1$ for some constant $C_{\varepsilon,\delta}$. Note that $\frac{g(x)}{x}=1-\varepsilon+\varepsilon x$, so there exists some $\alpha=\alpha_{\varepsilon,\delta}\in(0,1)$ such that for all $x\leq\alpha$, we have $g(x)\leq(1-\delta)x$. Thus it suffices to show there exists some nonnegative integer $N\leq C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)$ such that $g^{N}(\rho)\leq\alpha$. If $\rho\leq\alpha$, we simply take $N=0$. Otherwise, note that for $\alpha\leq x<1$, we have
$\frac{1-g(x)}{1-x}=1+\varepsilon-\varepsilon(1-x)\geq 1+\varepsilon\alpha>1,$
so we may take $N\geq\log_{1+\varepsilon\alpha}\left(\frac{1-\alpha}{1-\rho}\right).$
Remark 2.2.
We remark that this bound is essentially tight. Suppose all component functions defining our hierarchical multilinear function $f$ have non-linearity exactly $\varepsilon>0$, i.e., each component function $g$ satisfies $d(g,\operatorname{Lin})=\varepsilon$, and $\operatorname{Corr}(X_{i},Y_{i})=\rho$ for all $i$. With these tight conditions, following the proof of Lemma 1.1, it is straightforward to see that we could have proven $\operatorname{Corr}(f(X),f(Y))\geq(1-\varepsilon)\rho$ instead of (1.1). Inductively applying this bound instead of Lemma 1.1, we see that our hierarchical multilinear function of depth $d$ has $\operatorname{Corr}(f(X),f(Y))\geq(1-\varepsilon)^{d}\rho$. Hence, it is impossible to have exponential rate smaller than $1-\varepsilon$, and we have provided a bound with exponential rate $\alpha$ for all $\alpha>1-\varepsilon$.
Remark 2.3.
Note that in the case $\varepsilon=1$, i.e., $f$ is a hierarchical multilinear function where each component function is completely uncorrelated with any linear function, Lemma 1.1 actually implies $\operatorname{Corr}(f(X),f(Y))\leq\rho^{2^{d}}$. In this case, we have doubly exponential decay in depth, with decay rate depending on $\rho$, stronger than the singly exponential decay provided by Theorem 1.3, where the decay rate depends on $\varepsilon$.
Remark 2.4.
In the literature of cryptography, a Boolean function is $t$-resilient if it is balanced and remains unpredictable even if an adversary knows up to $t$ of the $n$ input bits. Concretely, this means $\widehat{f}(S)=0$ for all $\left\lvert S\right\rvert\leq t$. We refer interested readers to [7, Chapter 8] for further background on Boolean functions in cryptography. Our results do not care about the balanced condition, i.e., $\widehat{f}(\emptyset)=0$, but aside from this, our case $\varepsilon=1$ in the Boolean setting corresponds to 1-resilient functions. It is straightforward to see, by adapting the proof of Lemma 1.1, that for any positive integer $t$, if we have a hierarchical function whose components are all $t$-resilient Boolean functions, then we have $\operatorname{Corr}(f(X),f(Y))\leq\rho^{(t+1)^{d}}$.
Remark 2.5.
One example of such a component function is the parity function in the Boolean case, i.e., $\operatorname{Parity}(x)=x_{1}\cdot x_{2}\cdots x_{n}$, where $X_{i}\sim\operatorname{Unif}(\left\{\pm 1\right\})$ for all $i$. This function takes value 1 if an even number of bits are on, i.e., have value 1-1, and value 1-1 if an odd number of bits are on. In the coding theory literature, this function is more commonly written as
$\operatorname{Parity}(x_{1},\dots,x_{n})=x_{1}\oplus\cdots\oplus x_{n},$
where $\oplus$ denotes the XOR (exclusive OR), and each $x_{i}\in\{0,1\}$ is a bit. However, to fit this into our multilinear polynomial framework, we convert the additive group $\{0,1\}$ modulo 2, i.e., $\mathbb{Z}/2\mathbb{Z}$, into the multiplicative group $\{1,-1\}$, which is a standard transformation in the literature of analysis of Boolean functions. Assuming $\operatorname{Corr}(X_{i},Y_{i})=\rho$ for all $i$, it is well-known, and can be seen by adapting the proof of Lemma 1.1, that $\operatorname{Corr}(\operatorname{Parity}(X),\operatorname{Parity}(Y))=\rho^{n}$, i.e., exponential in $n$. Note that one can view the parity function as a hierarchical function of smaller component parity functions, and by expressing the $n$-bit parity function as a hierarchical function of depth $\lfloor\log_{2}n\rfloor$ with components being 2-bit parity functions, we have, via the $\varepsilon=1$ version of Theorem 1.3 discussed in Remark 2.3, that
$\operatorname{Corr}(\operatorname{Parity}(X),\operatorname{Parity}(Y))\leq\rho^{2^{\lfloor\log_{2}n\rfloor}}<\rho^{n/2},$
so our result still demonstrates exponential decay in $n$, albeit with a different rate than the true rate.
For example, note that Theorem 1.3 shows that the noise stability of the Boolean function on $3^{d}$ bits consisting of recursively composing the majority function on 3 bits, which is not linear, decays exponentially in $d$. Thus, the recursive majority function is noise sensitive for logarithmic depth. Recall that the fact that the recursive majority function is noise sensitive is folklore; see, for example, [4, 5].
Theorem 1.3 implies low-degree bounds on hierarchical multilinear functions. To make this concrete, the following lemma shows how bounds on noise stability, that is, $\operatorname{Corr}(f(X),f(Y))$, imply bounds on the maximum correlation between $f$ and any multilinear function of degree at most $D$.
Lemma 2.6.
Suppose $\{(X_{i},Y_{i}):i\in[n]\}$ are mutually independent, where for some $0\leq\rho\leq 1$, we have $\operatorname{Corr}(X_{i},Y_{i})=\rho$ for all $i$. Suppose $X_{i}$ and $Y_{i}$ have the same first and second moments, i.e., $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]$ and $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]<\infty$, for all $i\in[n]$. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a multilinear function, and let $M$ denote the maximum correlation between $f$ and any multilinear function of degree at most $D$, i.e.,
$M=\sup_{\begin{subarray}{c}g:\mathbb{R}^{n}\to\mathbb{R}\text{ multilinear}\\ \operatorname{deg}(g)\leq D\end{subarray}}\operatorname{Corr}(f(X),g(X)).$
Then
$M\leq\rho^{-D/2}\sqrt{\operatorname{Corr}(f(X),f(Y))}.$
Proof.
We wish to bound
$M=\sup_{\begin{subarray}{c}g:\mathbb{R}^{n}\to\mathbb{R}\text{ multilinear}\\ \operatorname{deg}(g)\leq D\end{subarray}}\operatorname{Corr}(f(X),g(X)).$
For any such $g$, decompose $f$ and $g$ as
$f=\sum_{S\subseteq[n]}\widehat{f}(S)\chi_{S},\quad g=\sum_{S\subseteq[n]}\widehat{g}(S)\chi_{S},$
and without loss of generality rescale $g$ so that $\operatorname{Var}(g)=1$, i.e.,
$\sum_{S\neq\emptyset}\widehat{g}(S)^{2}=1.$
Following the proof of Lemma 1.1, we know
$\operatorname{Corr}(f(X),f(Y))=\frac{1}{\operatorname{Var}(f)}\sum_{S\neq\emptyset}\rho^{\left\lvert S\right\rvert}\widehat{f}(S)^{2}.$
Similarly, i.e., via orthogonality of $\left\{\chi_{S}:S\subseteq[n]\right\}$, by the Cauchy–Schwarz inequality, we have
$\operatorname{Corr}(f(X),g(X))=\frac{1}{\sqrt{\operatorname{Var}(f)}}\sum_{S\neq\emptyset}\widehat{f}(S)\widehat{g}(S)=\frac{1}{\sqrt{\operatorname{Var}(f)}}\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)\widehat{g}(S)\leq\frac{\sqrt{\displaystyle\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}}}{\sqrt{\operatorname{Var}(f)}},$
with equality if and only if
$g=C+\frac{\displaystyle\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)\chi_{S}}{\sqrt{\displaystyle\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}}}$
for some constant $C\in\mathbb{R}$. Thus,
$M=\frac{1}{\sqrt{\operatorname{Var}(f)}}\sqrt{\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}}.$
Then we relate $\operatorname{Corr}(f(X),f(Y))$ with $M$ via
$\operatorname{Corr}(f(X),f(Y))$ $=\frac{1}{\operatorname{Var}(f)}\sum_{S\neq\emptyset}\rho^{\left\lvert S\right\rvert}\widehat{f}(S)^{2}$
$\geq\frac{1}{\operatorname{Var}(f)}\sum_{1\leq\left\lvert S\right\rvert\leq D}\rho^{\left\lvert S\right\rvert}\widehat{f}(S)^{2}$
$\geq\frac{1}{\operatorname{Var}(f)}\sum_{1\leq\left\lvert S\right\rvert\leq D}\rho^{D}\widehat{f}(S)^{2}$
$=\rho^{D}M^{2},$
which rearranges to our desired inequality. ∎
Using Theorem 1.3, for any hierarchical multilinear function $f$ of non-linearity $\varepsilon>0$ and depth $d$, and for any $\delta<\varepsilon$ and $0\leq\rho<1$, applying Lemma 2.6 yields (1.3). Note that this is because, as previously discussed in Remark 1.2, we can always construct a $\rho$-correlated $Y$ satisfying the conditions of Theorem 1.3. In particular, for fixed $D$, our bound (1.3) decays exponentially in depth $d$.
This in turn lets us prove Theorem 1.9.
Proof of Theorem 1.9.
From the proof of Lemma 2.6, we have for any $D\in[n]$,
$M=\frac{1}{\sqrt{\operatorname{Var}(f)}}\sqrt{\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}}\geq\sqrt{\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}}$
as $f(x)\in[-1,1]$ for all $x$. Then by (1.3), we have, fixing $\rho\in(0,1)$ to be a constant,
$\sqrt{\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}}\leq C_{\rho,\varepsilon,\delta}\rho^{-D/2}(1-\delta)^{d/2}=e^{\Theta(D)-\Theta(\log n)},$
for some constant $C_{\rho,\varepsilon,\delta}>0$. Pick constants $c_{1},c_{2}>0$ sufficiently small with $c_{2}\geq c_{1}$ so that, letting $D=\frac{c_{1}\log n}{\log\log n}$ and $x=n^{c_{2}}$, we have
$xe^{D}C_{\rho,\varepsilon,\delta}\rho^{-D/2}(1-\delta)^{d/2}=o(1).$
Let this expression be denoted by $\tau$. [8, Theorem 3.2] implies there exists constant $K>0$ such that $f$ is $O\left(\tau+\exp\left(-Kx^{2/D}\right)n^{2D+2}\right)$-approximately $D$-resilient. Note that, re-picking $c_{1}$ as necessary, this asymptotic expression is $o(1)$, because
$Kx^{2/D}+(2D+2)\log n=-K(\log n)^{\frac{2c_{2}}{c_{1}}}+\Theta\left(\frac{\log^{2}(n)}{\log\log(n)}\right)=-\omega(1).$
Applying [8, Theorem 1.1] then completes the proof. ∎
A similar proof yields Theorem 1.10.
Proof of Theorem 1.10.
Garban, Pete, and Schramm [11] showed that
$\sum_{1\leq\left\lvert S\right\rvert\leq D}\widehat{f}(S)^{2}=n^{-\frac{1}{2}+o(1)}D^{\frac{2}{3}+o(1)}$
for any $D\in\left[N^{3}\right]$. Let $D=\frac{c\log n}{\log\log n}$ and $x=n^{c}$ for some constant $c\in\left(0,\frac{1}{4}\right)$. Then we have
$xe^{D}\sqrt{\sum_{1\leq S\leq D}\widehat{f}(S)^{2}}=n^{c}n^{\frac{c}{\log\log n}}n^{-\frac{1}{4}+o(1)}=o(1)$
as $c<\frac{1}{4}$. Let this expression be denoted by $\tau$. Then [8, Theorem 3.2] implies there exists constant $K>0$ such that $f$ is $O\left(\tau+\exp\left(-Kx^{2/D}\right)N^{3(2D+2)}\right)$-approximately $D$-resilient. This asymptotic expression is $o(1)$ as $\tau=o(1)$ and
$Kx^{2/D}+(24D+24)\log n=-K(\log n)^{2}+\Theta\left(\frac{\log^{2}(n)}{\log\log(n)}\right)=-\omega(1).$
Applying [8, Theorem 1.1] then completes the proof. ∎
3. Non-separable functions
3.1. Correlated probability spaces
If we hope to generalize Theorem 1.3 beyond multilinear functions, we will need to strengthen our initial assumption, i.e., that $\operatorname{Corr}(X_{i},Y_{i})\leq\rho$. Because $\operatorname{Corr}(X_{i},Y_{i})$ only captures linear relationships between $X_{i}$ and $Y_{i}$, if our function is not multilinear, it has the potential to have higher correlation $\operatorname{Corr}(f(X),f(Y))$ than that of its components, i.e., $\operatorname{Corr}(X_{i},Y_{i})$, by utilizing these nonlinear relationships. Recall Examples 1.4 and 1.5, which illustrate how nonlinear relationships in a single variable can lead to noise stability that does not decay with depth. To this end, we broaden our concept of correlation via the following definition of correlation between probability spaces, sometimes referred to as the “Hirschfeld–Gebelein–Rényi maximal correlation” as it was first introduced by Hirschfeld [15] and Gebelein [12] and then studied by Rényi [25].
Definition 3.1.
Given a probability measure $\mathbb{P}$ defined on $\Omega=\prod_{i=1}^{k}\Omega_{i}$, we say that $\Omega_{1},\dots,\Omega_{k}$ are correlated spaces. Given two linear subspaces $A$ and $B$ of $L^{2}(\mathbb{P})$, we define the correlation between $A$ and $B$ by
$\rho(A,B;\mathbb{P})=\rho(A,B)=\sup\left\{\operatorname{Corr}(f,g):f\in A,g\in B\right\}$
For $i,j\in[k]$, we define the correlation $\rho(\Omega_{i},\Omega_{j};\mathbb{P})$ by
$\rho(\Omega_{i},\Omega_{j};\mathbb{P})=\rho(\Omega_{i},\Omega_{j})=\rho(L^{2}(\Omega_{i},\mathbb{P}),L^{2}(\Omega_{j},\mathbb{P});\mathbb{P}).$
As $\rho$ also denotes a parameter in this paper, we will also use the notation $\operatorname{Corr}(\cdot)$ in place of $\rho(\cdot)$ for clarity.
For random vector $X:\Omega\to\mathbb{R}^{n}$, let $L^{2}(X)$ denote the linear subspace of $L^{2}(\mathbb{P})$ that is $\sigma(X)$-measurable. Each $f\in L^{2}(X)$ can be identified with a function $\mathfrak{f}:\mathbb{R}^{n}\to\mathbb{R}$ satisfying $\mathbb{E}[\mathfrak{f}(X)^{2}]<\infty$, and we will frequently make no distinction between these two functions. Equipped with this new definition, we will consider the condition that $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))\leq\rho$, in place of the condition that $\operatorname{Corr}(X_{i},Y_{i})\leq\rho$.
We recall some theory from Mossel [21, §2.2] concerning noise correlation of functions. For two random variables $X,Y:\Omega\to\mathbb{R}$, if $f\in L^{2}(X)$ satisfies $\mathbb{E}[f(X)]=0$ and $\mathbb{E}[f(X)^{2}]=1$, then the maximizer of $\left\lvert\mathbb{E}[f(X)g(Y)]\right\rvert$ among $g\in L^{2}(Y)$ satisfying $\mathbb{E}[g(Y)^{2}]=1$ is given by
$g=\frac{\mathbb{E}[f(X)|Y]}{\left\lVert\mathbb{E}[f(X)|Y]\right\rVert_{2}}=\frac{\mathbb{E}[f(X)|Y]}{\sqrt{\mathbb{E}[\mathbb{E}[f(X)|Y]^{2}]}},$
and
$\left\lvert\mathbb{E}[f(X)g(Y)]\right\rvert=\left\lVert\mathbb{E}[f(X)|Y]\right\rVert_{2}$
The function $\mathbb{E}[f(X)|Y]\in L^{2}(Y)$ is the Markov operator from $L^{2}(X)$ to $L^{2}(Y)$, and as such we will denote this $g$ by $Tf\in L^{2}(Y)$. We have $\mathbb{E}[Tf]=0$, so
(3.1) $\operatorname{Corr}(L^{2}(X),L^{2}(Y))=\sup\left\{\mathbb{E}[f(X)Tf(X)]:f\in L^{2}(X),\mathbb{E}[f]=0,\mathbb{E}[f^{2}]=1\right\}.$
3.2. Efron–Stein decomposition
Returning to our setup in Section 2, where $\mathbb{P}$ is a product measure on product space $\Omega$ over the laws of correlated pairs of random variables $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, we now introduce the Efron–Stein decomposition as our analog to the $\chi_{S}$ that decomposed our multilinear polynomials.
Definition 3.2.
Let $(\Omega_{1},\mu_{1}),\dots,(\Omega_{n},\mu_{n})$ be laws of correlated pairs of real-valued random variables $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, and let $(\Omega,\mu)=\prod_{i=1}^{n}(\Omega_{i},\mu_{i})$. The Efron–Stein decomposition of $f\in L^{2}(X)$ is given by
$f(x)=\sum_{S\subseteq[n]}f_{S}(x),$
where each function $f_{S}$ depends only on $x_{S}$ and for all $S\not\subseteq S^{\prime}$, we have $\mathbb{E}[f_{S}|X_{S^{\prime}}]=0$ almost surely. We can similarly define the Efron–Stein decomposition for any $g\in L^{2}(Y)$.
It is well-known that the Efron–Stein decomposition exists and is unique [10]. A standard result (see, for example, [21, Proposition 2.11], which only proves the result in the discrete case but naturally holds in general) states that the Efron–Stein decomposition “commutes” with the Markov operator, i.e., $(Tf)_{S}=T(f_{S})\in L^{2}(Y)$, where $T:L^{2}(X)\to L^{2}(Y)$ is the Markov operator, which is given by $T=\bigotimes_{i=1}^{n}T_{i}$ for Markov operators $T_{i}:L^{2}(X_{i})\to L^{2}(Y_{i})$. It thus is unambiguous to write $Tf_{S}$. Similarly generalizing the proof of [21, Proposition 2.12], if $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))\leq\rho_{i}$ for some $\rho_{i}$ for each $i$, then for all $S\subseteq[n]$,
(3.2) $\left\lVert Tf_{S}\right\rVert_{2}\leq\left(\prod_{i\in S}\rho_{i}\right)\left\lVert f_{S}\right\rVert_{2}.$
3.3. Hierarchical non-separable functions
Recall the goal of this section is to generalize Theorem 1.3 beyond multilinear functions. Just as we had to broaden our assumption that $\operatorname{Corr}(X_{i},Y_{i})\leq\rho$ to $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))\leq\rho$, we must also broaden our assumption that $f$ is non-linear to account for non-linear relationships in a single variable. To this end, we replace the concept of non-linearity with non-separability, where recall an (additively) separable function $f:\mathbb{R}^{n}\to\mathbb{R}$ is a function of the form $f(x)=f_{1}(x_{1})+\cdots+f_{n}(x_{n})$ for $f_{1},\dots,f_{n}:\mathbb{R}\to\mathbb{R}$. Note that in the context of multilinear polynomials, separability and linearity are equivalent, so this definition of a separable function is consistent with our setup in Section 2. We can then define non-separability by the condition that
$\operatorname{Corr}(L^{2}(f(X)),L^{2}(X_{1})+\cdots+L^{2}(X_{n}))\leq 1-\varepsilon.$
For any $g\in L^{2}(f(X))$ and $h_{i}\in L^{2}(X_{i})$ for all $i\in[n]$, where we normalize $\mathbb{E}[g]=0=\mathbb{E}[h_{i}]$ for all $i$ and $\mathbb{E}[g^{2}]=1=\mathbb{E}[(h_{1}+\cdots+h_{n})^{2}]$, we have by the defining properties of the Efron–Stein decomposition, where we view $G=g\circ f\in L^{2}(X)$,
$\operatorname{Corr}(g,h_{1}+\cdots+h_{n})$ $=\mathbb{E}[g(f(X))(h_{1}(X_{1})+\cdots+h_{n}(X_{n}))]$
$=\sum_{S\neq\emptyset}\mathbb{E}[G_{S}(h_{1}+\cdots+h_{n})_{S}]$
$=\sum_{i=1}^{n}\mathbb{E}[G_{\{i\}}h_{i}]$
$\leq\sum_{i=1}^{n}\left\lVert G_{\{i\}}\right\rVert_{2}\left\lVert h_{i}\right\rVert$
$\leq\sqrt{\sum_{i=1}^{n}\left\lVert G_{\{i\}}\right\rVert_{2}^{2}}\sqrt{\sum_{i=1}^{n}\left\lVert h_{i}\right\rVert_{2}^{2}}$
$=\sqrt{\sum_{i=1}^{n}\left\lVert G_{\{i\}}\right\rVert_{2}^{2}}.$
Note that by picking $h_{i}$ appropriately, one can make both applications of the Cauchy–Schwarz inequality tight, and thus
(3.3) $\operatorname{Corr}(L^{2}(f(X)),L^{2}(X_{1})+\cdots+L^{2}(X_{n}))=\sup_{\begin{subarray}{c}g\in L^{2}(f(X))\\ \mathbb{E}[g]=0,\mathbb{E}[g^{2}]=1\end{subarray}}\sqrt{\sum_{i=1}^{n}\left\lVert G_{\{i\}}\right\rVert_{2}^{2}}.$
Because of the square root on the right hand side, we will reparametrize $\varepsilon$ so that our non-separability condition is
$\operatorname{Corr}(L^{2}(f(X)),L^{2}(X_{1})+\cdots+L^{2}(X_{n}))\leq\sqrt{1-\varepsilon}.$
By naturally extending our definition of $d(f,V)$ to $d(U,V)$ for $U,V\subseteq L^{2}(\Omega,\mathbb{P})$ by
$d(U,V)=\inf_{f\in U,g\in V}\frac{\mathbb{E}[(f-g)^{2}]}{\operatorname{Var}(f)},$
note that
$d(L^{2}(f(X)),L^{2}(X_{1})+\cdots+L^{2}(X_{n}))$ $=\inf_{g,h_{1},\dots,h_{n}}\frac{\displaystyle\sum_{S}\left\lVert G_{S}-(h_{1})_{S}-\cdots-(h_{n})_{S})\right\rVert_{2}^{2}}{\displaystyle\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}^{2}}$
(3.4) $=\inf_{\begin{subarray}{c}g\in L^{2}(f(X))\\ \mathbb{E}[g]=0,\mathbb{E}[g^{2}]=1\end{subarray}}\displaystyle\sum_{\left\lvert S\right\rvert\geq 2}\left\lVert G_{S}\right\rVert_{2}^{2},$
where we have used the fact that $f\mapsto f_{S}$ is a linear operator, which follows from linearity of expectation. Assuming $\mathbb{E}[g]=0$ and $\mathbb{E}[g^{2}]=1$, we have $G_{\emptyset}\equiv 0$ and $\sum_{S\subseteq[n]}\left\lVert G_{S}\right\rVert_{2}^{2}=1$, so we see that the extremizers in (3.3) and (3.4) are the same $g$, and letting $V=L^{2}(X_{1})+\cdots+L^{2}(X_{n})$,
(3.5) $\operatorname{Corr}(L^{2}(f(X)),V)\leq\sqrt{1-\varepsilon}\iff d(L^{2}(f(X)),V)\geq\varepsilon.$
While the latter condition is more consistent with our condition in Section 2, we will use the former condition to consistently use the language of correlated probability spaces. Of course, this choice is purely stylistic, as these two conditions are equivalent. We say a function $f\in L^{2}(X)$ satisfying this condition is $\varepsilon$-non-separable. Likewise, we can define a hierarchical function $f\in L^{2}(X)$ of non-separability $\varepsilon>0$ and depth $d$ analogously to how we defined hierarchical multilinear functions.
We are now ready to prove Lemma 1.6, our generalization of Lemma 1.1 to non-separable functions.
Proof of Lemma 1.6.
Let $Z=f(X)$ and $W=f(Y)$. Pick arbitrary $g,h:\mathbb{R}\to\mathbb{R}$ such that $\mathbb{E}[g(Z)]=\mathbb{E}[h(W)]=0$ and $\mathbb{E}[g(Z)^{2}]=\mathbb{E}[h(W)^{2}]=1$. It suffices to show
$\mathbb{E}[g(Z)h(W)]\leq(1-\varepsilon)\rho+\varepsilon\rho^{2}.$
Let
$G=g\circ f\in L^{2}(X)\quad\text{and}\quad H=h\circ f\in L^{2}(Y).$
Note that for distinct $S,S^{\prime}\subseteq[n]$, we have $\mathbb{E}[G_{S}(X)H_{S^{\prime}}(Y)]=0$. This is because we either have $S\not\subseteq S^{\prime}$ or $S^{\prime}\not\subseteq S$, and in the former case, by the defining properties of the Efron–Stein decomposition and recalling the $(X_{i},Y_{i})$ are mutually independent across $i$,
$\mathbb{E}[G_{S}(X)H_{S^{\prime}}(Y)]$ $=\mathbb{E}[\mathbb{E}[G_{S}(X)H_{S^{\prime}}(Y)|(X_{S^{\prime}},Y_{S^{\prime}})]]$
$=\mathbb{E}[\mathbb{E}[G_{S}(X)|(X_{S^{\prime}},Y_{S^{\prime}})]H_{S^{\prime}}(Y)]$
$=\mathbb{E}[0\cdot H_{S^{\prime}}(Y)]$
$=0.$
Thus,
$\mathbb{E}[g(Z)h(W)]$ $=\sum_{S\neq\emptyset}\mathbb{E}[G_{S}(X)H_{S}(Y)]$
$=\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\mathbb{E}\left[\frac{G_{S}(X)}{\left\lVert G_{S}\right\rVert_{2}}\frac{H_{S}(Y)}{\left\lVert H_{S}\right\rVert_{2}}\right]$
$\leq\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\mathbb{E}\left[\frac{G_{S}(X)}{\left\lVert G_{S}\right\rVert_{2}}T\left(\frac{G_{S}}{\left\lVert G_{S}\right\rVert_{2}}\right)\right]$
$=\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\left\lVert TG_{S}\right\rVert_{2}$
$\leq\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\rho^{\left\lvert S\right\rvert}$,
where the first inequality uses the property that the Markov operator maximizes covariance as discussed in Section 3.1, and the second inequality follows from (3.2). Applying the Cauchy–Schwarz inequality and then our non-separability condition, we bound
$\mathbb{E}[g(Z)h(W)]$ $\leq\sqrt{\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}^{2}\rho^{\left\lvert S\right\rvert}}\sqrt{\sum_{S\neq\emptyset}\left\lVert H_{S}\right\rVert_{2}^{2}\rho^{\left\lvert S\right\rvert}}$
$\leq(1-\varepsilon)\rho+\varepsilon\rho^{2},$
which completes the proof. ∎
Theorem 1.8 follows from the same proof as that of Theorem 1.3 using Lemma 1.6 in place of Lemma 1.1.
Remark 3.3.
Again, we note that in the case $\varepsilon=1$, i.e., $f$ is a hierarchical function where each component function has zero correlation with any function of a single one of its component variables, Lemma 1.6 implies $\operatorname{Corr}(L^{2}(f(X)),L^{2}(f(Y)))\leq\rho^{2^{d}}$. So, analogously to Remark 2.3, in this case we have doubly exponential decay in depth with decay rate depending on $\rho$, stronger than the singly exponential decay provided by Theorem 1.8 with decay rate depending on $\varepsilon$.
Next, we state and prove the following general version of Lemma 2.6, which we use to justify (1.4).
Lemma 3.4.
Suppose $\{(X_{i},Y_{i}):i\in[n]\}$ are mutually independent, where for some $0\leq\rho\leq 1$, we have $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))=\rho$ for all $i$. Suppose $X$ and $Y$ have the same marginal distributions. If $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfies $\mathbb{E}[f(X)^{2}]<\infty$, then let $M$ denote the maximum correlation between $L^{2}(f)$ and any function of Efron–Stein degree at most $D$, i.e., $M=\operatorname{Corr}(L^{2}(f(X)),\mathcal{G}),$ where $\mathcal{G}=\sum_{\left\lvert S\right\rvert\leq D}L^{2}(X_{S}).$ Let
$\gamma=\sup_{G,H}\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\rho^{\left\lvert S\right\rvert}$,
where the supremum is taken over all $g,h:\mathbb{R}\to\mathbb{R}$ such that $\mathbb{E}[g(f(X))]=\mathbb{E}[h(f(Y))]=0$ and $\mathbb{E}[g(f(X))^{2}]=\mathbb{E}[h(f(Y))^{2}]=1$, and $G=g\circ f\in L^{2}(X)$ and $H=h\circ f\in L^{2}(Y)$. Then
$M\leq\rho^{-D/2}\sqrt{\gamma}.$
Proof.
Let $Z=f(X)$ and $W=f(Y)$. Using a slightly generalized but essentially similar argument as that which we used to show (3.3), we have
$M=\sup_{\begin{subarray}{c}g\in L^{2}(f(X))\\ \mathbb{E}[g]=0,\mathbb{E}[g^{2}]=1\end{subarray}}\sqrt{\sum_{1\leq\left\lvert S\right\rvert\leq D}\left\lVert G_{S}\right\rVert_{2}^{2}},$
where $G=g\circ f\in L^{2}(X)$. It then suffices to show that for any such $g$,
$\sqrt{\sum_{1\leq\left\lvert S\right\rvert\leq D}\left\lVert G_{S}\right\rVert_{2}^{2}}\leq\rho^{-D/2}\left(\sup_{\begin{subarray}{c}h\in L^{2}(f(Y))\\ \mathbb{E}[h]=0,\mathbb{E}[h^{2}]=1\end{subarray}}\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\rho^{\left\lvert S\right\rvert}\right)^{1/2},$
where $H=h\circ f\in L^{2}(Y)$. Letting $g$ and $h$ be the same when viewed as functions $\mathbb{R}\to\mathbb{R}$, because $X$ and $Y$ have the same marginal distributions, we note that $H_{S}=G_{S}$ for all $S$, which achieves the equality case of the Cauchy–Schwarz inequality that implies
$\sup_{\begin{subarray}{c}h\in L^{2}(f(Y))\\ \mathbb{E}[h]=0,\mathbb{E}[h^{2}]=1\end{subarray}}\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\rho^{\left\lvert S\right\rvert}=\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}^{2}\rho^{\left\lvert S\right\rvert}$.
Hence, it suffices to show
$\sum_{1\leq\left\lvert S\right\rvert\leq D}\left\lVert G_{S}\right\rVert_{2}^{2}\rho^{D}\leq\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}^{2}\rho^{\left\lvert S\right\rvert}$,
which is clearly true as $0\leq\rho\leq 1$. ∎
To justify (1.4), note that for any $0\leq\rho<1$, by Remark 1.7, we can construct $Y$ satisfying the conditions of Lemma 3.4. Recall that Theorem 1.8 was proven by inductively applying Lemma 1.6. Using the fact that the proof of Lemma 1.6 implies the stronger technical statement
$\sup_{G,H}\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\rho^{\left\lvert S\right\rvert}\leq(1-\varepsilon)\rho+\varepsilon\rho^{2},$
the proof of Theorem 1.8 implies that for a hierarchical function $f\in L^{2}(X)$ of non-separability $\varepsilon>0$ and depth $d$, and for any $\delta<\varepsilon$ and $0\leq\rho<1$ (where because our constructed $Y$ has the same marginal distribution as $X$, we have $f$ is also such a hierarchical function in $L^{2}(Y)$),
$\sup_{G,H}\sum_{S\neq\emptyset}\left\lVert G_{S}\right\rVert_{2}\left\lVert H_{S}\right\rVert_{2}\rho^{\left\lvert S\right\rvert}\leq(1-\delta)^{d-C_{\varepsilon,\delta}\log\left(\frac{1}{1-\rho}\right)}.$
Thus, Lemma 3.4 implies (1.4).
Lastly, we use Theorem 1.8 to prove Theorem 1.11.
Proof of Theorem 1.11.
By [9, Theorem 1.7], it suffices to show that for any polynomial $p:\mathbb{R}^{m}\to\mathbb{R}$ of degree less than $D=n^{c}$, we have
$\mathbb{E}[\left\lvert f(X)-p(X)\right\rvert]\geq\frac{1}{8}-o(1).$
To show this we will use [9, Theorem 1.5]. By Theorem 1.8,
$\operatorname{Corr}(L^{2}(f(X)),L^{2}(f(Y)))\leq(1-\delta)^{d-\Theta(\log\frac{1}{1-\rho})}.$
Fixing $\delta<\varepsilon$ and setting $1-\rho=\frac{1}{D^{3}}$, this bound is $e^{-\Theta(\log n)+\Theta(\log D)}$, and choosing $c$ sufficiently small ensures this is $o(1)$. For any such polynomial $p$, by [9, Theorem 1.5], we have
$\mathbb{E}[\left\lvert f(X)-p(X)\right\rvert]\geq\frac{\mathbb{P}(f(X)\neq f(Y))}{4}-O(D\sqrt{1-\rho})=\frac{\frac{1}{2}-\frac{1}{2}\mathbb{E}[f(X)f(Y)]}{4}-o(1),$
where $Y$ is a $\rho$-correlated spherical Gaussian. By (3.1), $\operatorname{Corr}(L^{2}(X_{i}),L^{2}(Y_{i}))=\rho$ for all $i$, so
$\mathbb{E}[\left\lvert f(X)-p(X)\right\rvert$] $\geq\frac{1}{8}-\frac{\mathbb{E}[f(X)f(Y)]}{8}-o(1)$
$\geq\frac{1}{8}-\frac{\operatorname{Var}(f(X))\operatorname{Corr}(L^{2}(f(X)),L^{2}(f(Y)))}{8}-o(1)$
$\geq\frac{1}{8}-\frac{(1-\delta)^{d-\Theta(\log\frac{1}{1-\rho})}}{8}-o(1)$
$=\frac{1}{8}-o(1),$
which completes the proof. ∎
Acknowledgements
E.M. was partially supported by NSF DMS-2031883, Bush Faculty Fellowship ONR-N00014-20-1-2826, and Simons Investigator award (622132).